# Kinematic-ICP: Enhancing LiDAR Odometry with Kinematic Constraints for Wheeled Mobile Robots Moving on Planar Surfaces

### v1

@@ Abstract
LiDAR odometry is essential for many robotics applications, including 3D mapping, navigation, and simultaneous localization and mapping. LiDAR odometry systems are usually based on some form of point cloud registration to compute the ego-motion of a mobile robot. Yet, few of today’s LiDAR odometry systems consider the domain-specific knowledge and the kinematic model of the mobile platform during the point cloud alignment. In this paper, we present Kinematic-ICP, a LiDAR odometry system that focuses on wheeled mobile robots equipped with a 3D LiDAR and moving on a planar surface, which is a common assumption for warehouses, offices, hospitals, etc. Our approach introduces kinematic constraints within the optimization of a traditional point-to-point iterative closest point scheme. In this way, the resulting motion follows the kinematic constraints of the platform, effectively exploiting the robot’s wheel odometry and the 3D LiDAR observations. We dynamically adjust the influence of LiDAR measurements and wheel odometry in our optimization scheme, allowing the system to handle degenerate scenarios such as feature-poor corridors. We evaluate our approach on robots operating in large-scale warehouse environments, but also outdoors. The experiments show that our approach achieves top performances and is more accurate than wheel odometry and common LiDAR odometry systems. Kinematic-ICP has been recently deployed in the Dexory fleet of robots operating in warehouses worldwide at their customers’ sites, showing that our method can run in the real world alongside a complete navigation stack.
@@ II Related Work
There are many practical solutions proposed for LiDAR odometry in the context of mobile robotics [29, 6, 7, 10, 22, 18, 35, 5]. Most of these methods build upon the ICP algorithm [2, 3] to align consecutive LiDAR point clouds and estimate the ego-motion of the sensor. Although some algorithms include proprioceptive sensors to improve the estimation [33, 32, 34], LiDAR odometry approaches typically do not systematically consider the kinematics of wheeled mobile robots. Our work aims to change this.
@@ V-A1 Datasets
The second platform we use to evaluate our approach is a Clearpath Husky A200, which is a four-wheeled with skid steering. The robot provides a wheel odometry estimate using its wheel encoders and is equipped with a Hesai LiDAR XT32, which records data at 10 Hz. We drive and record data on a wide and flat pavement on our campus and in a park in front of the Poppelsdorf palace in Bonn, which consists of uneven and rough terrain such as grass and curb walks. Note that this second location is challenging because it does not strictly follow our assumption of operating on a flat surface. To evaluate the accuracy of our approach, we mount a reflective prism on the robot and track the prism with a Leica Nova MS60 total station as shown in Fig. 2. The total station and the robot are initially time synchronized, and the prism can be tracked with an angular accuracy of 1 ” and a distance accuracy of 2 mm. This type of reference system has been previously used to accurately track the pose of a robot [27, 28, 26].
@@ V-A2 Baselines
We evaluate our method against the original version of KISS-ICP [29] and with two variants that incorporate the robot’s wheel odometry as an initial estimate for point cloud registration. In the first variant, we estimate the full 6-degree-of-freedom pose in 3D space, referred to as WO + 3D KISS. In the second variant, the motion is constrained to a 2D plane, simplifying the estimation process. This is denoted as WO + 2D KISS.
In addition to comparing our method with various KISS-ICP [29] variants, we further evaluate its performance by benchmarking it against two widely used state estimation methods in the robotics industry [14]. Specifically, we focus on integrating the most effective KISS-ICP [29] variant, WO + 2D KISS, with the Fuse framework[14] and the robot_localization package [16]. The Fuse framework is a fixed-lag smoother that optimizes a pose graph by incorporating wheel odometry, LiDAR odometry, and kinematic constraints to estimate the robot’s position. Meanwhile, robot_localization is an extended Kalman filter (EKF) that fuses sensor data to estimate the robot’s state.
@@ V-A3 Metrics
To evaluate the different methods, we use the average translation error from the KITTI Odometry [8] benchmark, which calculates the average relative error over various trajectory lengths in percent. This metric is commonly used in autonomous driving, where the segment lengths for evaluation are considerably large for a wheeled mobile robot. To address this, we use smaller interval lengths of 1, 2, 5, 10, 20, 50, and 100 meters. Additionally, we report the root mean squared absolute translation error (ATE) after alignment as a measure of the global drift in the estimated trajectories.
@@ V-B Qualitative Results in Large Indoor Warehouses
We first show on the top left the raw wheel odometry in Fig. 3, which follows the kinematic model of a differential drive but does drift over time. Next, we compare the result with two modified versions of KISS-ICP to account for the planar motion, and we observe that the racks become more visible, but the trajectories still bend. Note that the publicly available implementation of KISS-ICP will fail on this sequence due to the sparseness of the sensor [19]. Besides that, the estimated trajectory is not smooth, which is unreasonable for a 12 meter tall robot that weighs approximately 500 kg. Finally, using our proposed approach, the estimation is globally more accurate and still smooth. This is a direct consequence of enforcing the kinematic constraints in the optimization, which guarantees mathematically that the produced pose is always smooth and locally consistent.
@@ V-C Quantitative Results
The EKF and Fuse baselines perform well by combining LiDAR and wheel odometry, with Fuse often ranking second. However, our approach consistently outperforms both in terms of ATE and RPE by directly processing the data and incorporating a kinematic model. Additionally, as noted earlier, our method is more computationally efficient since it optimizes wheel odometry and LiDAR corrections directly rather than fusing two different sources of odometry. For comparison, our system runs at Hz on a single-core CPU, while Fuse runs at approximately Hz.

### v2

@@ Abstract
LiDAR odometry is essential for many robotics applications, including 3D mapping, navigation, and simultaneous localization and mapping. LiDAR odometry systems are usually based on some form of point cloud registration to compute the ego-motion of a mobile robot. Yet, few of today’s LiDAR odometry systems consider domain-specific knowledge or the kinematic model of the mobile platform during the point cloud alignment. In this paper, we present Kinematic-ICP, a LiDAR odometry system that focuses on wheeled mobile robots equipped with a 3D LiDAR and moving on a planar surface, which is a common assumption for warehouses, offices, hospitals, etc. Our approach introduces kinematic constraints within the optimization of a traditional point-to-point iterative closest point scheme. In this way, the resulting motion follows the kinematic constraints of the platform, effectively exploiting the robot’s wheel odometry and the 3D LiDAR observations. We dynamically adjust the influence of LiDAR measurements and wheel odometry in our optimization scheme, allowing the system to handle degenerate scenarios such as feature-poor corridors. We evaluate our approach on robots operating in large-scale warehouse environments, but also outdoors. The experiments show that our approach achieves top performances and is more accurate than wheel odometry and common LiDAR odometry systems. Kinematic-ICP has been recently deployed in the Dexory fleet of robots operating in warehouses worldwide at their customers’ sites, showing that our method can run in the real world alongside a complete navigation stack.
@@ II Related Work
There are many practical solutions proposed for LiDAR odometry in the context of mobile robotics [29, 6, 7, 10, 22, 18, 35, 5]. Most of these methods build upon the ICP algorithm [2, 3] to align consecutive LiDAR point clouds and estimate the ego-motion of the sensor. Although some algorithms include proprioceptive sensors to improve the estimation [33, 32, 34], LiDAR odometry approaches often do not systematically consider the kinematics of wheeled mobile robots. Our work aims to change this.
@@ V-A1 Datasets
The second platform we use to evaluate our approach is a Clearpath Husky A200, which is a four-wheeled with skid steering. The robot provides a wheel odometry estimate using its wheel encoders and is equipped with a Hesai LiDAR XT32, which records data at 10 Hz. We drive and record data on a wide and flat pavement on our campus and in a park in front of the Poppelsdorf Palace in Bonn, which consists of uneven and rough terrain such as grass and curb walks. Note that this second location is challenging because it does not strictly follow our assumption of operating on a flat surface. To evaluate the accuracy of our approach, we mount a reflective prism on the robot and track the prism with a Leica Nova MS60 total station as shown in Fig. 2. The total station and the robot are initially time-synchronized. The prism can be tracked with an angular accuracy of 0.0003 ∘ and a range accuracy of 3 mm or better. This type of reference system has been previously used to accurately track the pose of a robot [27, 28, 26].
@@ V-A2 Baselines
We evaluate our method against the original version of KISS-ICP [29] and with two variants that incorporate the robot’s wheel odometry as an initial estimate for point cloud registration. In the first variant, we estimate the full 6-degree-of-freedom pose in 3D space, referred to as WO + 3D KISS-ICP. In the second variant, the motion is constrained to a 2D plane, simplifying the estimation process. This is denoted as WO + 2D KISS-ICP.
In addition to comparing our method with various KISS-ICP [29] variants, we further evaluate its performance by benchmarking it against two widely used state estimation methods in the robotics industry [14]. Specifically, we focus on integrating the most effective KISS-ICP [29] variant, WO + 2D KISS-ICP, with the Fuse framework[14] and the robot_localization package [16]. The Fuse framework is a fixed-lag smoother that optimizes a pose graph by incorporating wheel odometry, LiDAR odometry, and kinematic constraints to estimate the robot’s position. Meanwhile, robot_localization is an extended Kalman filter (EKF) that fuses sensor data to estimate the robot’s state.
@@ V-A3 Metrics
To evaluate the different methods, we use the average translation error from the KITTI Odometry [8] benchmark, which calculates the average relative error over various trajectory lengths in percent. This metric is commonly used in autonomous driving, where the segment lengths for evaluation are considerably large for a wheeled mobile robot. To address this, we use smaller interval lengths of 1, 2, 5, 10, 20, 50, and 100 m. Additionally, we report the root mean squared absolute translation error (ATE) after alignment as a measure of the global drift in the estimated trajectories.
@@ V-B Qualitative Results in Large Indoor Warehouses
We first show on the top left the raw wheel odometry in Fig. 3, which follows the kinematic model of a differential drive but does drift over time. Next, we compare the result with two modified versions of KISS-ICP to account for the planar motion, and we observe that the racks become more visible, but the trajectories still bend. Note that the publicly available implementation of KISS-ICP will fail on this sequence due to the sparseness of the sensor [19]. Besides that, the estimated trajectory is not smooth, which is unreasonable for a 12 m tall robot that weighs approximately 500 kg. Finally, using our proposed approach, the estimation is globally more accurate and still smooth. This is a direct consequence of enforcing the kinematic constraints in the optimization, which guarantees mathematically that the produced pose is always smooth and locally consistent.
@@ V-C Quantitative Results
The EKF and Fuse baselines perform well by combining LiDAR and wheel odometry, with Fuse often ranking second. However, our approach consistently outperforms both in terms of ATE and RPE by directly processing the data and incorporating a kinematic model. Additionally, as noted earlier, our method is more computationally efficient since it optimizes wheel odometry and LiDAR corrections directly rather than fusing two different sources of odometry. For comparison, our system runs at 100 Hz on a single-core CPU, while Fuse runs at approximately Hz.